solving rational expressions and equations

7 over 8 + 2 over 5 = x over 20
show work

Answer by fractalier(6550)

The first step in solving almost all fractional equations is to multiply all the terms by the lowest common denominator...here that is 40...so we get
7/8 + 2/5 = x/20
40(7/8 + 2/5 = x/20)
35 + 16 = 2x
2x = 51
x = 51/2